please help me the question is a take home test.
in chemistry class, 9 liters of a 4% SILVER IODIDE solution must be mixed with a 10% solution to get a 6% solution. how many liters of the 10% solution are needed?

Answer by stanbon(75887)

in chemistry class, 9 liters of a 4% SILVER IODIDE solution must be mixed with a 10% solution to get a 6% solution. how many liters of the 10% solution are needed?
-----------------
4% solution DATA;
Amount = 9 liter ; amt of active ingredient = 0.04*9= 0.36 liter
---------------
10% solution DATA:
Amount = x liter ; amt of active ingredient = 0.10x
-----------------
6% Mixture DATA:
Amount = 9+x liter; amt of active ingredient = 0.06(9+x)=0.54+0.06x liter
======================
EQUATION:
active + active = active
0.36 + 0.10x = 0.54 + 0.06x
0.04x = 0.18
x = 4.5 liter (amt of 10% solution that must be added to get the mixture)
==================
Cheers,
Stan H.